OpenStudy (anonymous):

If 4 is the determinant of matrix A, what is arcdet(A)?

OpenStudy (anonymous):

What is arcdet?

OpenStudy (anonymous):

The determinant of arcA

OpenStudy (anonymous):

what is arc A? The inverse?

OpenStudy (zzr0ck3r):

surprisingly 1/4

OpenStudy (zzr0ck3r):

people use arcA to talk about matrix inverse?

OpenStudy (anonymous):

Ok so will it always be 1/A?

OpenStudy (zzr0ck3r):

the determinant of the inverse is the inverse of the determinant.

OpenStudy (zzr0ck3r):

1/det(A)

OpenStudy (anonymous):

Thanks zzr0ckey

OpenStudy (zzr0ck3r):

fo shee z

OpenStudy (anonymous):

Yep. Since:\[AA^{-1}=I\]Taking determinants yields:\[\det(AA^{-1})=\det(I)\Longrightarrow \det(A)\det(A^{-1})=1\]\[\Longrightarrow \det(A^{-1})=\frac{1}{\det(A)}\]

OpenStudy (zzr0ck3r):

A*A^-1 =1 det(A8A^-1) = det(1)

OpenStudy (zzr0ck3r):

bah* your to fast:)

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):